The coil is rotated through an angle of 90° in 2.0 second. Calculate the magnitude of the average e.m.f. induced in the coil while it is being rotated.


1
Expert's answer
2021-05-26T10:33:50-0400

"E=|\frac{d\Phi}{dt}|=\frac{\Delta\Phi}{\Delta t}=\frac{1}{2}\Delta\Phi"


Assume that "\Delta\Phi=8\ (Wb)". Then


"E=\frac{1}{2}\cdot8=4\ (V)" . Answer